Amy, Bobby, and Candace are teaming up to paint a wall. If all three work together, it would take 80 minutes. If Amy and Candace work together, it would take 2 hours. If only Bobby and Candace work together, it would take 144 minutes. Suppose Amy and Bobby work together for 1 hour, then leave. If Candace arrives and must finish painting the wall by her self, how much longer will it take her to finish painting the wall?
Answer by richwmiller(17219)

1/a+1/b+1/c=1/80,
1/a +1/c=1/120,
1/b+1/c=1/144,
1/120+1/b=1/80,
8/960+1/b=12/960
1/b=4/960
1/b=1/240
b=240
1/240+1/c=1/144
3/720+1/c=5/720
1/c=2/720,
c=360
1/a+1/360=1/120
1/a+1/360=3/360
1/a=2360=1/180
a=180= 3 hours
b=240= 4 hours
c=360= 6 hours
Suppose Amy and Bobby work together for 1 hour, then leave. If Candace arrives and must finish painting the wall by her self, how much longer will it take her to finish painting the wall?
60/180+60/240+x/c=1
1/3+1/4+x/360=1
4/12+3/12+x/360=1
7/12+x/360=1
x/360=5/12
x/360=150/360
x=150 minutes for Candace to finish 2.5 hours or 2 hours and 30 minutes